Find the equation of the straight line which
passes through the point(-3,5) and is
perpendicular to the line 2x-4y+3=0?

Answer by greenestamps(13334)



Using basic algebra....

(1) Find the slope of the given line

2x-4y+3=0
4y = 2x+3
y = (1/2)x+3/4

The slope is 1/2; the slope of a line perpendicular to that line is -2/1, or just -2.

(2) Use the general slope-intercept form of the equation with the desired slope of -2 passing through (x,y) = (-3,5) to find the equation

y = mx+b
5 = (-2)(-3)+b
5 = 6+b
b = -1

ANSWER: y = -2x-1

A faster path to the answer....

Given the equation 2x-4y+3=0, any line parallel to that line has the form 2x-4y+C=0 for some constant C, and any line perpendicular to that line has the form 4x+2y+C=0 for some constant C (switch the coefficients of x and y and change the sign of one of them).

So we are looking for an equation of the form 4x+2y+C=0 that is satisfied by (x,y) = (-3,5):

4(-3)+2(5)+C=0
-12+10+C=0
C=2

The equation we want is

4x+2y+2=0

That is equivalent to the answer we got using basic algebra:

4x+2y+2=0
2y = -4x-2
y = -2x-1